4.0os um hikness, what is the area of the (D)
Answer #1

density d = 19.32 g/cm^3 = 19.32*10^3 kg/m^3

mass m = 3.006*10^-3 kg

a) volume v = area*thickness

v = A*t

density = mass/volume

volume = mass/density

A*t = 3.006*10^-3/19.32*10^3 = 0.155*10^-6

A = 0.155*10^-6/4.005*10^-6

A = 0.0388 m^2

2) volume of cylindrical fiber is v = pi*r^2*l

pi*r^2*l = mass/density = 0.155*10^-6

l = 0.155*10^-6/3.14*2.0^2*10^-12

l = 12381.47 m

Please recheck the input values, some were not visible properly.